OpenStudy (rainbow_dashie):

For which operations is the set {0, 1} closed? Choose all answers that are correct.

OpenStudy (rainbow_dashie):

A. multiplication B. division C. addition D. subtraction

OpenStudy (anonymous):

C

OpenStudy (rainbow_dashie):

thx Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct.

OpenStudy (rainbow_dashie):

A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (rainbow_dashie):

a. (-1) d.(-1)

OpenStudy (rainbow_dashie):

@bibby

OpenStudy (rainbow_dashie):

@ganeshie8